2. What is the molar concentration of a AgNO3 solution if 16.35-mL react with:
a) 0.3017-g KIO3 ANSWER: 0.08623 M AgNO3
b) 12.33 mL of 0.02149 M K4Fe(CN)6 ANSWER: 0.06482 M AgNO3
Molar masses: KIO3 = 214.0 K4Fe(CN)6 = 368.4
1
Expert's answer
2020-06-04T10:16:10-0400

Solution:

a)

The balanced chemical equation:

AgNO3 + KIO3 = AgIO3 + KNO3

According to the chemical equation: n(AgNO3) = n(KIO3)


Moles of KIO3 = n(KIO3) = Mass of KIO3 / Molar mass of KIO3

n(KIO3) = (0.3017 g) / (214.0 g/mol) = 0.00141 mol

n(AgNO3) = n(KIO3) = 0.00141 mol


Molarity of AgNO3 = Moles of AgNO3 / Volume of solution

Molarity of AgNO3 = (0.00141 mol) / (0.01635 L) = 0.08624 mol/L = 0.08624 M

Molarity of AgNO3 is 0.08624 M.


Answer (a): The molar concentration of a AgNO3 solution is 0.08624 M.


b)

The balanced chemical equation:

4AgNO3 + K4[Fe(CN)6] = Ag4[Fe(CN)6] + KNO3

According to the chemical equation: n(AgNO3)/4 = n(K4[Fe(CN)6])


Moles of K4[Fe(CN)6] = n(K4[Fe(CN)6]) = Molarity of K4[Fe(CN)6] × Volume of solution

n(K4[Fe(CN)6]) = (0.02149 M) × (0.01233 L) = 0.000265 mol

n(AgNO3) = 4 × n(K4[Fe(CN)6]) = 4 × (0.000265 mol) = 0.00106 mol


Molarity of AgNO3 = Moles of AgNO3 / Volume of solution

Molarity of AgNO3 = (0.00106 mol) / (0.01635 L) = 0.06483 mol/L = 0.06483 M

Molarity of AgNO3 is 0.06483 M.


Answer (b): The molar concentration of a AgNO3 solution is 0.06483 M.